Assuming the complete dissociation of HCl and the lead salt ,calculate how much HCl is to be added to a 0.001 M lead salt solution to just prevent precipitation when saturated with  H2S?The concentration of H2S in its saturated solution is 0.1 M. KaH2S=1.1×10-23;KspPbS=3.4×10-28.

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Ankur,

[Pb+2] [S2-] = Ksp = 3.4 * 10^-28

[PbS] = [Pb+2] = 0.001   (saturated)

[S2-]= 3.4 * 10^-25

Ka = 1.1 * 10^-23 = [H+]2[S-2]/[H2S]

[H+]2 = 1.1/3.4 * 10^2 * 0.1

[H+] = 1.8

since due to the very small Ka value of H2S the H+ furnished by it is negligible.

Therefore the 1.8 M HCl is to be added.